OpenStudy (anonymous):

i have 2 questions please help meee Solve the system of equations. x-4y=5 x=3y-2 (___,___) Which numbers belong to the range of the relation {(–2, 3), (5, –1), (0, 3), (5, –2)}? Choose all answers that are correct. A. –2 B. –1 C. 0 D. 3

OpenStudy (haseeb96):

Do u know how to solve the first one just isolate the x from the equation no#1 and then put the value of that isolated term into the other equation at the place of x

OpenStudy (anonymous):

@Haseeb96 i dont get it

OpenStudy (anonymous):

@Haseeb96 hello o.o

OpenStudy (haseeb96):

x-4y=5.......equation #1 x=3y-2,,,,,,,equation# 2 put x =3y-2 at the place of x in the equation #1 it would be 3y-2 -4y =5 -y=5+2 y=-7 now put y=-7 into the equation # 2 x = 3(-7) -2 = -21 -2 x = -23 so (x,y) = (-23,-7) this is the correct answer

OpenStudy (anonymous):

oh my god thankyou thankyou thankyou

OpenStudy (anonymous):

dude ily XD

OpenStudy (anonymous):

can you also help me with the second one, ive been stuck on it forever

OpenStudy (anonymous):

sorry im annoying XD

OpenStudy (haseeb96):

-2,-1 and 3 are the correct answer